# Limitations of tensor network approaches for optimization and sampling: A comparison against quantum and classical Ising machines

### v1

@@ Abstract
Optimization problems pose challenges across various fields. In recent years, quantum annealers have emerged as a promising platform for tackling such challenges. To provide a new perspective, we develop a heuristic tensor-network-based algorithm to reveal the low-energy spectrum of Ising spin-glass systems with interaction graphs relevant to present-day quantum annealers. Our deterministic approach combines a branch-and-bound search strategy with an approximate calculation of marginals via tensor-network contractions. Its application to quasi-two-dimensional lattices with large unit cells of up to 24 spins, realized in current quantum annealing processors, requires a dedicated approach that utilizes sparse structures in the tensor network representation and GPU hardware acceleration. We benchmark our approach on random problems defined on Pegasus and Zephyr graphs with up to a few thousand spins, comparing it against the D-Wave Advantage quantum annealer and Simulated Bifurcation algorithm, with the latter representing an emerging class of classical Ising solvers. Apart from the quality of the best solutions, we compare the diversity of low-energy states sampled by all the solvers. For the biggest considered problems with over 5000 spins, the state-of-the-art tensor network approaches lead to solutions that are $0.1\%$ to $1\%$ worse than the best solutions obtained by Ising machines while being two orders of magnitude slower. We attribute those results to approximate contraction failures. While all three methods can output diverse low-energy solutions, e.g., differing by at least a quarter of spins with energy error below $1\%$, our deterministic branch-and-bound approach finds sets of a few such states at most. On the other hand, both Ising machines prove capable of sampling sets of thousands of such solutions.
@@ I Introduction
Many physics-inspired approaches have been developed to tackle that problem. A prominent and general family of methods, starting with the Simulated Annealing algorithm [4], employs thermal fluctuations introduced by Metropolis-Hastings updates [5, 6] within a Markov-chain Monte Carlo (MCMC) paradigm. Rugged energy landscapes pose significant challenges for local Monte-Carlo updates, motivating elaborate schemes to navigate distinct local minima. Those include Parallel Tempering (PT) [7] to enhance thermal fluctuations, or heuristic cluster moves, like Iso-energetic or Houdayer cluster move [8, 9], or Nonequilibrium Monte Carlo with inhomogeneous temperature distribution over identified surrogate problem backbones [10].
@@ V.1 Problem instances
The probability distribution defining instances in class III was selected for Pegasus geometry through a trial-and-error method to ensure that problems from this class are challenging for classical algorithms. Specifically, Ref. [49] presents extensive benchmarks comparing the time-to-solution between D-Wave Advantage quantum annealer and a selection of classical algorithms, including simulated annealing, parallel tempering with isoenergetic cluster moves, integer quadratic programming, message passing, etc. (but not SBM). It demonstrated that in this class, and for some range of target approximation ratios, quantum annealer provides run-time benefits over a tested collection of established classical methods.
@@ V.2 Metrics: best energy, diversity of solutions
In the first metric, we focus on the lowest energy obtained with given resources, which we quantify using solver runtime. We will be plotting a relative energy difference above the best energy found, $d_{E}=(E-E_{\mathrm{best}})/2|E_{\mathrm{best}}|$, where the reference energy $E_{\mathrm{best}}$ is established by collecting the results of all tested solvers. We use the same definition to specify approximation ratio $a_{r}$, which refers to all solutions with $d_{E}\leq a_{r}$.
In the second metric, we probe the quality of sampling of diverse approximate solutions of a spin-glass problem [50, 51]. We summarize the idea behind the diversity measure below. The goal is to select—from the set of available low-energy solutions with energies within a given approximation ratio $a_{r}$—the largest set of independent solutions. The size of the set provides a diversity of solutions. We refer to two spin configurations ${\underline{s}}_{N}$ and ${\underline{s^{\prime}}}_{N}$, using here original spin variables, as independent if their Hamming distance $d$ is above a given threshold
@@ V.3 Results of the ground state search
Figure 10 shows the time to median approximation ratio for 20 instances of class I for Pegasus geometry with 216 spins, 1176 spins, and 5376 spins and Zephyr graph with 332 spins and 563 spins. Analogous data for other instance classes (with similar performance) are collected in the Appendix. Apart from full-space TN simulations, we additionally conduct simulations with local dimensional reduction of cluster degrees of freedom, as described in Sec. IV. Here, we cut the $2^{24}$ cluster states to $2^{20}$ and $2^{16}$ states for Pegasus instances, and from $2^{16}$ states to $2^{14}$ and $2^{12}$ states for Zephyr instances. Due to the memory limitations, for Pegasus instances with 5376 spins, we only performed calculations with local dimensional reduction to $2^{16}$ states. The parameter selection was performed separately for each type of graph: an inverse temperature $\beta=0.5$ was used for Pegasus instances, while $\beta=1.0$ was used for Zephyr instances. We further show the influence of $\beta$ on the quality of the results in the following section and in Appendix C.
@@ Appendix A Auxiliary results
In Fig. A1, we show time to approximation ratio for class II and III instances achieved by all examined solvers: D-Wave QA, SBM and TN approach. Time to diversity ratio for classes I, II and III with the distance threshold $R=1/2$ and $R=1/4$ are shown in the Fig. A2 and A3, respectively.
Figure A1: Time to approximation ratio. Here, we show the results for instances in class II and III for Pegasus and Zephyr geometries for a selection of system sizes. The data complement the results in Fig. 10 and were obtained with the same simulation parameters.
Figure A2: Time to diversity ratio for $R=1/2$. We show the time needed to reach a given fraction of diverse solutions within $a_{r}=0.01$ and $R=1/2$ for instance class II and III, complementing the data in Fig. 11.
Figure A3: Time to diversity ratio for $R=1/4$. We show the data for all three instance classes and approximation ratio $a_{r}=0.01$, complementing similar results in Figs. 11 and A2 for $R=1/2$. For large system sizes, the fraction of diverse solutions identified by TN becomes negligible as compared to two other solvers, where a limited available width of a branch and bound procedure $M=1024$, apparently, allows identifying only a small fraction of independent solutions for $\mathcal{D}_{\rm total}$ in hundreds. The simulation parameters are the same as in Fig. 10.

### v3

# Limitations of tensor-network approaches for optimization and sampling: A comparison to quantum and classical Ising machines

@@ Abstract
Optimization problems pose challenges across various fields. In recent years, quantum annealers have emerged as a promising platform for tackling such challenges. To provide a new perspective, we develop a heuristic tensor-network-based algorithm to reveal the low-energy spectrum of Ising spin-glass systems with interaction graphs relevant to present-day quantum annealers. Our deterministic approach combines a branch-and-bound search strategy with an approximate calculation of marginals via tensor-network contractions. Its application to quasi-two-dimensional lattices with large unit cells of up to 24 spins, realized in current quantum annealing processors, requires a dedicated approach that utilizes sparse structures in the tensor network representation and GPU hardware acceleration. We benchmark our approach on random problems defined on Pegasus and Zephyr graphs with up to a few thousand spins, comparing it against the D-Wave Advantage quantum annealer and Simulated Bifurcation algorithm, with the latter representing an emerging class of classical Ising solvers. Apart from the quality of the best solutions, we compare the diversity of low-energy states sampled by all the solvers. For the biggest considered i.i.d. problems with over 5000 spins, the state-of-the-art tensor network approaches lead to solutions that are $0.1\%$ to $1\%$ worse than the best solutions obtained by Ising machines while being two orders of magnitude slower. We attribute those results to approximate contraction failures. For embedded tile planting instances, our approach gets to approximately $0.1\%$ from the planted ground state, a factor of $3$ better than the Ising solvers. While all three methods can output diverse low-energy solutions, e.g., differing by at least a quarter of spins with energy error below $1\%$, our deterministic branch-and-bound approach finds sets of a few such states at most. On the other hand, both Ising machines prove capable of sampling sets of thousands of such solutions.
@@ I Introduction
Many physics-inspired approaches have been developed to tackle that problem. A prominent and general family of methods, starting with the Simulated Annealing algorithm [4], employs thermal fluctuations introduced by Metropolis-Hastings updates [5, 6] within a Markov-chain Monte Carlo (MCMC) paradigm. Rugged energy landscapes pose significant challenges for local Monte-Carlo updates, motivating elaborate schemes to navigate distinct local minima. Those include Parallel Tempering (PT) [7] to enhance thermal fluctuations, or heuristic cluster moves, like Iso-energetic or Houdayer cluster move [8, 9], or Nonequilibrium Monte Carlo with inhomogeneous temperature distribution over identified surrogate problem backbones [10]. Alternative approaches to overcoming limitations of local updates include subgraph-based methods, which leverage fixed-treewidth substructures to improve sampling efficiency [11]. Additionally, highly optimized implementations of simulated annealing for Ising spin glasses have been demonstrated to achieve significant performance gains through multi-spin coding, particularly targeting previous graph geometry of D-Wave quantum annelers, i.e., Chimera graphs [12].
@@ IV Local dimensional reduction
We note that within the temperature range considered in this study, BP messages do not fully converge with the maximal change of messages during iteration saturating at around $10^{-3}$. Hence, reaching saturation in convergence, messages after a predefined number of iterations are used to select representatives of each cluster. At sufficiently high temperatures (small $\beta$), BP would be expected to converge. However, such temperature regimes are not practically relevant to the optimization problems studied here. Finally, we should note that the selection based on local marginals may not be able to identify local configurations contributing to a globally optimal state.
@@ V.1 Problem instances
  • class IV: We utilize tile planting instances on a square spin lattice, following the method outlined in Ref. [52, 53]. The instances defined on a square lattices with the system sizes $L=10$, $L=20$, $L=40$ are subsequently embedded into Pegasus graphs. After embedding, the number of spins increases to 164, 721 and 3065 spins respectively. The instances are constructed with a planted solution, meaning that the ground state is known a priori. The complexity of the planted instances is controlled by the probabilities $p_{1}$, $p_{2}$ and $p_{3}$, such that $p_{1}+p_{2}+p_{3}\leq 1$. We set $p_{2}=1.0$ for which the instances are expected to be the hardest to solve.
@@
The probability distribution defining instances in class III was selected for Pegasus geometry through a trial-and-error method to ensure that problems from this class are challenging for classical algorithms. Specifically, Ref. [51] presents extensive benchmarks comparing the time-to-solution between D-Wave Advantage quantum annealer and a selection of classical algorithms, including simulated annealing, parallel tempering with isoenergetic cluster moves, integer quadratic programming, message passing, etc. (but not SBM). It demonstrated that in this class, and for some range of target approximation ratios, quantum annealer provides run-time benefits over a tested collection of established classical methods. A recent comparison with the SBM algorithm was performed in [54] on max-cut problems showing the competitive performance of SBM.
@@ V.2 Metrics: best energy, diversity of solutions
In the first metric, we focus on the lowest energy obtained with given resources, which we quantify using solver runtime. We will be plotting a relative energy difference above the best energy found,
$d_{E}=(E-E_{\mathrm{best}})/2|E_{\mathrm{best}}|,$ (14)
where the reference energy $E_{\mathrm{best}}$ is established by collecting the results of all tested solvers when the ground truth is unknown.
In the second metric, we probe the quality of sampling of diverse approximate solutions of a spin-glass problem [55, 56]. We summarize the idea behind the diversity measure below. We focus on good-quality solutions with energies within a given approximation ratio $a_{r}$, that is, the solutions for which $d_{E}\leq a_{r}$. The goal is to select from the available set of such solutions the largest set of independent solutions. The size of the set provides a diversity of solutions. We refer to two spin configurations ${\underline{s}}_{N}$ and ${\underline{s^{\prime}}}_{N}$, using here original spin variables, as independent if their Hamming distance $d$ is above a given threshold
@@ V.3 Results of the ground state search
Figure 10 shows the time to median approximation ratio for 20 instances of class I for Pegasus geometry with 216 spins, 1176 spins, and 5376 spins and Zephyr graph with 332 spins and 563 spins. Analogous data for other instance classes are collected in the Appendix, where we observe similar performance for instances in classes I-III. For the largest considered problems in class IV, the TN algorithm outperforms QA and SBM in terms of the best energy found, improving $d_{E}$ by a factor of $3$ to $4$.
Apart from full-space TN simulations, we additionally conduct simulations with local dimensional reduction of cluster degrees of freedom, as described in Sec. IV. Here, we cut the $2^{24}$ cluster states to $2^{20}$ and $2^{16}$ states for Pegasus instances, and from $2^{16}$ states to $2^{14}$ and $2^{12}$ states for Zephyr instances. Due to the memory limitations, for Pegasus instances with 5376 spins, we only performed calculations with local dimensional reduction to $2^{16}$ states. The parameter selection was performed separately for each type of graph: an inverse temperature $\beta=0.5$ was used for Pegasus instances, while $\beta=1.0$ was used for Zephyr instances. We further show the influence of $\beta$ on the quality of the results in the following section and in Appendix C.
@@ Appendix A Auxiliary results
In Fig. A1, we show time to approximation ratio for class II, III and IV instances achieved by all examined solvers: D-Wave QA, SBM and TN approach. In contrast to classes I, II, and III, which exhibit similar performance trends, class IV instances demonstrate an advantage in energy results achieved by the TN approach. Although TN outperforms both D-Wave QA and SBM in the terms of energy for instances with 721 and 3065 spins, it fails to identify their ground states.
Time to diversity ratio for classes I, II, III and IV with the distance threshold $R=1/2$ and $R=1/4$ are shown in the Fig. A2 and A3, respectively.
Figure A1: Time to approximation ratio. Here, we show the results for instances in class II, III and VI for Pegasus and Zephyr geometries for a selection of system sizes. The data complement the results in Fig. 10 and were obtained with the same simulation parameters. We show the median based on 20 instances.
Figure A2: Time to diversity ratio for $R=1/2$. We show the time needed to reach a given fraction of diverse solutions within $a_{r}=0.01$ and $R=1/2$ for instance class II, III and IV, complementing the data in Fig. 11.
Figure A3: Time to diversity ratio for $R=1/4$. We show the data for all four instance classes and approximation ratio $a_{r}=0.01$, complementing similar results in Figs. 11 and A2 for $R=1/2$. For large system sizes, the fraction of diverse solutions identified by TN becomes negligible as compared to two other solvers, where a limited available width of a branch and bound procedure $M=1024$, apparently, allows identifying only a small fraction of independent solutions for $\mathcal{D}_{\rm total}$ in hundreds. The simulation parameters are the same as in Fig. 10.